In a town, 30% of the men and 25% of the women are overweight. If the town has 55% men and 45% women, what percent of the people are overweight?
Answer by ikleyn(52775)

.
If the entire population is P individuals, then


    the number of men   is  0.55*P,  and the number of overweight men   is 0.30*0.55*P.

    the number of women is  0.45*P,  and the number of overweight women is 0.25*0.45*P.


In all, the number of overweight individuals is  

    0.30*0.55*P + 0.25*0.45*P = (0.30*0.55 + 0.25*0.45)*P = 0.2775*P,


which gives the percent of the overweight population  27.75%.

Solved.